# Graph RAG for Legal Norms: A Hierarchical and Temporal Approach

### v2

@@ Abstract
This article proposes an adaptation of Graph Retrieval Augmented Generation (Graph RAG) specifically designed for the analysis and comprehension of legal norms, which are characterized by their predefined hierarchical structure, extensive network of internal and external references and multiple temporal versions. By combining structured knowledge graphs with contextually enriched text segments, Graph RAG offers a promising solution to address the inherent complexity and vast volume of legal data. The integration of hierarchical structure and temporal evolution into knowledge graphs - along with the concept of comprehensive Text Units - facilitates the construction of richer, interconnected representations of legal knowledge. Through a detailed analysis of Graph RAG and its application to legal norm datasets, this article aims to advance the field of Artificial Intelligence applied to Law, creating opportunities for more effective systems in legal research, legislative analysis, and decision support.
Keywords: Retrieval-Augmented Generation; Knowledge Graphs; Legal Tech; Legal AI; Legal Information Retrieval; Knowledge Representation.
@@ 1 Introduction
Additionally, the dynamic nature of legal norms, evolving through amendments, repeals, and new interpretations, results in multiple temporal versions. The ability to track and understand this evolution [2] is, therefore, essential for determining the applicable legislation at a given moment.
@@
Graph RAG presents significant potential for the analysis of legal norms. This technique can integrate structured information from knowledge graphs with unstructured textual content to create richer and more interconnected representations of legal knowledge. This integration allows capturing the complex relationships between legal entities, such as laws, doctrine, jurisprudence, concepts, and the temporal evolution of legal texts. By representing these legal entities, and their relationships, in a graph and linking them to the actual text of the legal norms and their different versions, Graph RAG can provide a more holistic understanding of Law [5].
@@
This paper is structured as follows. Section 2 first defines the core structural entities – Norm, Component, and Version – essential for modeling legal documents, discussing their hierarchical relationships and the challenges of representing temporal changes. Section 3 then elaborates on our proposed adaptations to the Graph RAG framework, introducing the concept of Text Units, explaining the rationale for modeling versions as aggregations rather than simple compositions, detailing the inclusion of essential metadata and alteration actions, and discussing strategies for structure-aware retrieval. Lastly, Section 4 concludes the paper by highlighting the key contributions and suggesting avenues for future investigation.
@@ 2 Structural Entities in the Graph
2.1 Legal Norm Structure
To structure the text of a Brazilian legal norm [9], the following components are defined [^5] In a more precise representation of the structuring of a legal norm, there would be the element Componente Autônomo Articulado (Autonomous Articulated Component), which would group all parts of the norm, including annexes. But, for didactic simplification, it will not be presented.:
@@ 2.1 Legal Norm Structure
The basic unit of a Brazilian legal norm’s articulation is the Artigo (Article). Each Article can be further subdivided into additional provisions, as demonstrated in Figure 1.
@@
Articles can also be organized into Dispositivos de Agrupamento (Grouping Provisions), as illustrated in Figure 2.
@@
As an example, Figure 3 presents the provision “Art. 12” of the “Constituição Federal Brasileira de 1988” (Federal Constitution of Brazil (1988)), displaying its grouping provisions (“Título II” and “Capítulo III”) and some of its child provisions (or subdivisions). Highlighted is the article, the basic unit of articulation. Annotations in orange are representing the type of each provision.
Figure 3: Example of articulated text for Art. 12 of the Federal Constitution of Brazil (1988) with annotations indicating the types of hierarchical provisions
As proposed, there will be 3 (three) types of Structural Entities for building the Knowledge Graph:
  1. 1.
    Norm (as a Whole): represents a legal norm.
  2. 2.
    Component: represents an element of the hierarchical structure of a legal norm. It can optionally be composed of other components (its subdivisions).
  3. 3.
    Version: version of the content (e.g., label and text) of a component, or of a norm (as a whole), over time – will be detailed further ahead.
The Norm and Component types are found in the Hierarchy dimension. The Version type is in the Time dimension [10].
@@ 2.2 Legal Norm Text Segmentation
In the original Graph RAG proposal, the first step involves segmenting the entire document into smaller text segments, or “text chunks”. In the subsequent step, each segment is processed—typically using a large language model (LLM)—to extract the identified entities, along with their properties and relationships, which are then added to the Knowledge Graph. However, when dealing with legal texts, the hierarchical structure is a fundamental characteristic that must be preserved to ensure accurate interpretation. Therefore, it is essential to extract and represent this hierarchical structure within the Knowledge Graph by mapping it to the previously defined structural entities.
@@
As an example, a passage of the the “Constituição Federal Brasileira de 1988” (Federal Constitution of Brazil (1988)) received the semantic segmentation (like showed in Figure 3), producing the entity of type Norm, represented by a circle in green, and those of type Component, represented in blue in the hierarchy presented in Figure 4.
For a more precise representation in the Graph, specialized entity types, derived from Norm, can be created for each type of norm in the treated legal system. In Brazil, there are, for example: "Constituição Federal" (Federal Constitution), “Lei Ordinária" (Ordinary Law) etc. Thus, the entity representing a legal norm could be included in the graph using the specialized type.
Specialized entity types, derived from Component, can also be created for each type of provision found in the treated legal system. For example, in the Brazilian case, there are types like "Artigo" (Article), "Inciso” (Item), “Alínea" (Subitem), “Título” (Title), “Capítulo” (Chapter), presented in figures 1 and 2. Thus, the entity representing a component could be included in the graph using the specialized type.
Figure 4: Illustration of hierarchical semantic segmentation, and typification of the structural entities, applied to a passage of the Federal Constitution of Brazil (1988), representing the "Norm" (green), "Components" (blue) and their associated texts.
@@ 2.3 Text Versioning over Time
Version entities represent the view over time of a norm/component, which can be extracted from its original text, but also from the “compilação multivigente999“Compilação” (Compilation) in Brazil is a process in which, starting from the original version of a legal norm, all modifications commanded by subsequent norms are applied to its text, up to the specified date, thus producing an unofficial updated text of this norm on this date. It is ”Multivigente” (Multi-temporal) when all dates on which the norm underwent alteration are considered in the Compilation, producing an unofficial updated text (version) for each identified date. (multi-temporal compilation) of its text.
A component always has an original version, which corresponds to the text, and label, of the component on the signature date of the original norm. It optionally has versions created by subsequent norms that commanded, for example, alteration of the provision to new content (e.g., alteration of the wording of an article by an amendment), or commanded a repeal (e.g., an article being explicitly repealed by a later law), or a suppression (e.g., an item being suppressed by a later norm), or a renumbering (e.g., an article being renumbered, label altered, by a later law) etc. of that component.
Since the text of a norm is constructed by hierarchically grouping the texts of its individual components, the creation of a new version for any component automatically generates a new version for the entire norm as well as for all grouping elements/components that include that component.
For example, Figure 5 illustrates the Version entities for the provision “Art. 6” (and its caput - main section) of ’Constituição Federal brasileira’ (Brazilian Federal Constitution)" [^10] The structuring of the ”Constituição Federal Brasileira de 1988” (Brazilian Federal Constitution of 1988) into ”Norm”, ”Components”, and ”Versions” can be visualized at https://normas.leg.br/?urn=urn:lex:br:federal:constituicao:1988-10-05;1988. Accessed on 15.03.2025., demonstrating how different versions reflect changes over time.
Figure 5: Structure of the provision "Art. 6" (and its caput - main section) of Brazilian Federal Constitution, illustrating four Version entities over time (light gray circles) and their respective generated events and resulting contents. The versions represent the different wordings of the caput on different dates, due to constitutional amendments.
@@ 2.4 Building the Graph
After semantically segmenting the text of a legal norm into a hierarchical structure (tree), as per the proposal presented above, the elements of the hierarchy are then inserted as nodes of the types Norm (with its associated properties, such as URN - Uniform Resource Name, name, signature date etc.) and Component (with its properties such as URN, optional inclusion date etc.) into the Knowledge Graph, as illustrated in Figure 4, which demonstrates the semantic segmentation of a Constitution passage.
Next, for each component, a node of the type Version will be inserted with its associated text segment [11] (and its properties such as URN, version date, label/name, text etc.) as presented in Figure 6 – this associated text segment (text chunk) is fundamental for the Retrieval Augmented Generation process, as it represents the textual content of that specific version that will be retrieved and used as context for the responses generated by the LLM.
Figure 6: Representation in the knowledge graph of the relationship between hierarchical structure entities "Norm", "Component”, their original versions "Version", and the associated Text Chunks.
For example, the Norm node/entity “Constituição Federal de 1988” (Federal Constitution of 1988)
identified by urn [^11] The formation of the urns presented in this work follows the standard proposed by the Brazilian LexML project. Available at: https://projeto.lexml.gov.br. Accessed on: 15 Mar 2025.
has a Component node/entity “caput do Art. 6º” (caput is the main text of the article) identified by
which has associated the original version Version node/entity, identified by
that represents the version on the date 1988-10-05 (original version) of the text of the caput of Art. 6 of Federal Constitution of 1988. For this Version node/entity, we have a associated Text chunk with the text bellow - as illustrated in Figure 5:
@@
Unlike other fields, where the text remains static over time (for example, the transcribed speech of a Parliamentarian in a Congress session will not change over time), the text over time of a legal norm, or of a component, can be represented by a set of Versions.
In Figure 7, four versions of the caput of Art. 6, found in Figure 5, are displayed: one original and three resulting from changes (text change) promoted by “Emendas Constitucionais” (Constitutional Amendments) - for a more precise representation in the Graph, specialized entity types, derived from Version, can be created for each type of version found in the treated legal system. For example, in Brazil, there are types like "Original Version”, “Alteração de Texto” (Text Change Version), “Renumeração” (Renumbering Version), “Revogação” (Repeal Version) etc. Thus, the entity representing the version could be included in the graph using the specialized type.
Figure 7: Representation of the multiple "Version" entities for the caput do Art. 6° provision (…!art6_cpt), illustrating its textual evolution over time and the link of each version to its Text Chunk.
The proposal of this work is not only to incorporate the entities of the hierarchical structure of the legal norm’s text into the knowledge graph, but also to include the Version entities of each component of the norm. It is important to note that the text segments (text chunks), identified by the original text and by the versions of the text over time of the norm, will be linked to the Version entities, and not to Component entities, as they represent the content of the component at different moments in time, as shown in Figure 7 - in practice, this link could be materialized as a property of the Version entity in the graph, just like a property for the embedding of its content.
This temporal representation will allow answering questions about the text of a component considering the temporal context: for example, if there was any change in the text of a component over time; what changes the component underwent over time, or within a period; what was the text of the component on a specific date, etc.
@@ 3.1 Text Units
Regarding the text segment (text chunk) associated with each Version entity, in cases of subdivisions of a caput, such as “parágrafos” (paragraphs) and enumeration elements (“incisos” (items) and “alíneas” (subitems)), we suggest adopting the strategy proposed by [12]. In this strategy, the text segment to be considered for a component is not only the text directly resulting from the semantic segmentation of the norm (or from its compilation over time) but the semantically complete text of the component, derived from post-segmentation (and post-compilation) processing. The cited author recognizes the hierarchical and interdependent nature of the components of a legal norm, where, for example, the meaning of an enumeration component (e.g., “inciso” (item)) is frequently dependent on the context of its parent component (e.g., the caput).
In other words, when including in the graph, for example, the original Version of the provision "inciso I do caput do Art. 5º” (item I of caput of Art. 5), presented in Figure 4, the text linked to this entity will not be just:
@@
but rather the text formed by the content of the article’s caput (on the same date) plus the content of the item itself, which can also be visualized in Figure 8 - the text of the caput was placed as a prefix in square brackets.
@@
Figure 8: Example of contextual (cumulative) Text Units for "incisos’ (items) of caput of Art. 5, demonstrating the inclusion of the parental caput text to provide complete semantic context.
The version of an article’s caput will have as linked text its own text accumulated with the text from the tree of its child components, as exemplified in Figure 9. The same applies to the article, the groupers and the norm (as a whole), accumulating its own text with the text from their respective trees.
According to the cited author, this would bring more “aboutness” [^12] The term ”aboutness” refers to the capacity of a text to faithfully represent its central theme. For theoretical discussions, see [17]. to the texts linked to the entities [13]. In other words, the cumulative “Text Unit”, by including the textual context of the parent, or child, components, represents the central theme and the legal meaning of the component in question more completely and comprehensively, increasing how well the text captures "what it is about" the provision.
Figure 9: Example of cumulative Text Unit for the caput of Art. 5 of the "Constituição Federal" (Federal Constitution).
Considering this new approach, we will call the texts associated with the entities Text Units, instead of Text Chunks, as they are not necessarily the pieces of text resulting directly from the segmentation (or compilation) of the norms’ texts, but rather the contextually complete content of each component.
3.2 Versions as Aggregations
In addition to the original version of a component C of a legal norm N, for each date/time T on which a subsequent norm Y commands an alteration in the content of C, a new version T must be produced for component C, as presented in Figure 10.
Figure 10: Versions of component C over time
Considering the proposal of Text Units resulting from the accumulation of text in the hierarchy, consequently, a new version T must also be produced for every component of norm N that:

### v3

# Graph RAG for Legal Norms: A Hierarchical, Temporal and Deterministic Approach

@@ Abstract
This article proposes an adaptation of Graph Retrieval-Augmented Generation (Graph RAG) specifically designed for the analysis and comprehension of legal norms. Legal texts are characterized by a predefined hierarchical structure, an extensive network of references and a continuous evolution through multiple temporal versions. This temporal dynamism poses a significant challenge for standard AI systems, demanding a deterministic representation of the law at any given point in time. To address this, our approach grounds the knowledge graph construction in a formal, FRBRoo-inspired model that distinguishes abstract legal works from their concrete textual expressions. We introduce a multi-layered representation of Temporal Versions (capturing date-specific changes) and Language Versions (capturing linguistic variations). By modeling normative evolution as a precise sequence of these versioned entities, we enable the construction of a knowledge graph that serves as a verifiable "ground truth". This allows Large Language Models to generate responses based on accurate, context-aware, and point-in-time correct legal information, overcoming the risk of temporal inaccuracies. Through a detailed analysis of this formal Graph RAG approach and its application to legal norm datasets, this article aims to advance the field of Artificial Intelligence applied to Law, creating opportunities for more effective and reliable systems in legal research, legislative analysis, and decision support.
Keywords: Retrieval-Augmented Generation; Knowledge Graphs; Legal Tech; Legal AI; Legal Information Retrieval; Temporal Modeling; Knowledge Representation.
@@ 1 Introduction
Additionally, the dynamic nature of legal norms, evolving through amendments, repeals, and new interpretations, results in multiple temporal versions. The ability to track and understand this evolution [2] is, therefore, essential for determining the applicable legislation at a given moment. This temporal dynamism presents a critical challenge for current AI systems, which, being inherently probabilistic, often cannot deterministically reconstruct the precise text of a law as it existed on a specific date. To ensure verifiable accuracy, our proposal is grounded in a formal, component-level versioning model, as detailed in [3], which provides a robust theoretical foundation for representing the diachronic evolution of legal texts.
@@
Graph RAG presents significant potential for the analysis of legal norms. This technique can integrate structured information from knowledge graphs with unstructured textual content to create richer and more interconnected representations of legal knowledge. This integration allows capturing the complex relationships between legal entities, such as laws, doctrine, jurisprudence, concepts, and the temporal evolution of legal texts. By representing these legal entities, and their relationships, in a graph and linking them to the actual text of the legal norms and their different versions, modeled as a precise sequence of temporal expressions, Graph RAG can provide a more holistic understanding of Law [7].
@@
This paper is structured as follows. Section 2 first defines the core structural entities – Norm, Component, and Version – and introduces the formal model for their representation, which is essential for modeling legal documents, discussing their hierarchical relationships and the challenges of representing temporal changes. Section 3 then elaborates on our proposed adaptations to the Graph RAG framework, introducing the concept of Text Units, explaining the rationale for modeling temporal versions as aggregations rather than simple compositions, detailing the inclusion of essential metadata and alteration actions, and discussing strategies for structure-aware retrieval. Lastly, Section 4 concludes the paper by highlighting the key contributions and suggesting avenues for future investigation.
@@ 2 Structural Entities in the Graph
2.1 A Formal Model for Structural and Temporal Representation
To precisely model the structural and temporal complexity of legal norms, we ground our approach in the Functional Requirements for Bibliographic Records (FRBRoo) framework [11], adapted for the legal domain as detailed in [3]. This model distinguishes between abstract concepts and their concrete textual realizations, enabling a deterministic representation of legal text evolution. We define four primary types of structural entities:
  1. 1.
    Norm (as a Whole): Represents a legal norm as an abstract intellectual creation (e.g., the "1988 Brazilian Federal Constitution"). This corresponds to the Work (F1 Work) concept in FRBRoo, capturing the legal instrument’s identity independent of any specific text or amendment.
  2. 2.
    Component: Represents a hierarchical element within a norm (e.g., a title, chapter, or article) as an abstract concept. Each component is a Component Work, an identifiable part of the main Work that persists conceptually even as its text changes over time.
  3. 3.
    Temporal Version (TV / CTV): Represents the precise text of a Norm or Component as it exists at a specific point in time. A Component Temporal Version (CTV) is the TV of a specific Component. Each is a distinct, date-stamped entity, corresponding to the FRBRoo Expression (F2 Expression) concept.
  4. 4.
    Language Version (LV / CLV): Represents the text of a specific Temporal Version in a particular language. A Component Language Version (CLV) is a LV of a specific CTV. Each is also an Expression that derives from a single TV.
For the sake of clarity and to avoid terminological clutter, we will adopt a simplified notation for the remainder of this paper. The concepts that apply to a Norm as a whole (e.g., Temporal Version) and those that apply to its parts (e.g., Component Temporal Version) are structurally identical. Therefore:
  • We will use the term Temporal Version to refer to both the temporal version of a Norm and a Component.
  • We will use the term Language Version to refer to both the language version of a Norm and a Component.
The context will make it clear whether the entity refers to the norm as a whole or to one of its specific components.
The Norm and Component types exist in the conceptual Hierarchy dimension. The Temporal Version introduces the Time dimension [12], while the Language Version adds a Linguistic dimension.
For the sake of simplicity in the subsequent examples and diagrams, which focus on the Brazilian monolingual official context, we will often refer to a single Version entity. This entity should be understood as the primary Temporal Version and its corresponding Portuguese Language Version combined. This pragmatic simplification allows us to focus on the hierarchical and temporal modeling, while acknowledging that the underlying framework fully supports multilingual representations.
2.2 Legal Norm Structure in Practice
To structure the text of a Brazilian legal norm [13], the following components are defined [^5] In a more precise representation of the structuring of a legal norm, there would be the element Componente Autônomo Articulado (Autonomous Articulated Component), which would group all parts of the norm, including annexes. But, for didactic simplification, it will not be presented.. Each of these structural elements, such as Artigo or Capítulo, is modeled as a Component entity in our graph.
@@ 2.2 Legal Norm Structure in Practice
The basic unit of a Brazilian legal norm’s articulation is the Artigo (Article). Each Article can be further subdivided into additional provisions (Components), as demonstrated in Figure 1.
@@
Articles can also be organized into Dispositivos de Agrupamento (Grouping Provisions/Components), as illustrated in Figure 2.
@@
As an example, Figure 3 presents the provision (Component) “Art. 12” of the “Constituição Federal Brasileira de 1988” (Federal Constitution of Brazil (1988)), displaying its grouping provisions/Components (“Título II” and “Capítulo III”) and some of its child provisions (Components). Highlighted is the article, the basic unit of articulation. Annotations in orange are representing the type of each provision (Component).
Figure 3: Example of articulated text for Art. 12 of the Federal Constitution of Brazil (1988) with annotations indicating the types of hierarchical provisions/components.
@@ 2.3 Legal Norm Text Segmentation
In the original Graph RAG proposal, the first step involves segmenting the entire document into smaller text segments, or “text chunks”. In the subsequent step, each segment is processed—typically using a large language model (LLM)—to extract the identified entities, along with their properties and relationships, which are then added to the Knowledge Graph. However, when dealing with legal texts, the hierarchical structure is a fundamental characteristic that must be preserved to ensure accurate interpretation. Therefore, it is essential to extract and represent this hierarchical structure within the Knowledge Graph by mapping it to the previously defined structural entities: Norm (as a Work), Component (as a Component Work), and their corresponding Versions (as Expressions).
@@
As an example, a passage of the the “Constituição Federal Brasileira de 1988” (Federal Constitution of Brazil (1988)) received the semantic segmentation (like showed in Figure 3), producing the abstract entity of type Norm, represented by a circle in green, and those of type Component, represented in blue in the hierarchy presented in Figure 4.
For a more precise representation in the Graph, specialized entity types, derived from Norm, can be created for each type of norm in the treated legal system. In Brazil, there are, for example: "Constituição" (Constitution), "Emenda Constitucional" (Constitutional Amendment), “Lei Ordinária" (Ordinary Law), "Lei Complementar" (Complementary Law) etc. Thus, the entity representing a legal Work could be included in the graph using the specialized type.
Specialized entity types, derived from Component, can also be created for each type of provision found in the treated legal system. For example, in the Brazilian case, there are types like "Artigo" (Article), "Inciso” (Item), “Alínea" (Subitem), “Título” (Title), “Capítulo” (Chapter), presented in figures 1 and 2. Thus, the entity representing a Component Work could be included in the graph using the specialized type.
Figure 4: Illustration of hierarchical semantic segmentation, and typification of the structural entities, applied to a passage of the Federal Constitution of Brazil (1988), representing the Norm (Work) in green, Components (Component Works) in blue, and their associated "texts chunks".
@@ 2.4 Text Versioning over Time
As defined in our formal model, the textual evolution of a norm or component is captured through Temporal Version (TV / CTV) entities. Each TV / CTV represents the precise state of a text at a specific point in time and can be extracted from its original enactment or from a “compilação multivigente999“Compilação” (Compilation) in Brazil is a process in which, starting from the original version of a legal norm, all modifications commanded by subsequent norms are applied to its text, up to the specified date, thus producing an unofficial updated text of this norm on this date. It is ”Multivigente” (Multi-temporal) when all dates on which the norm underwent alteration are considered in the Compilation, producing an unofficial updated text (version) for each identified date. (multi-temporal compilation). Each TV / CTV, in turn, has at least one Language Version (LV / CTV) containing the text in a specific language.
A Component always has an original Temporal Version (CTV), which corresponds to the text (and label) enacted on the signature date of the original norm. Subsequent legislative acts (e.g., amendments) that alter, repeal, suppress, or renumber a component do not change the abstract Component (the Component Work) itself, but rather create a new Temporal Version for it. This new CTV captures the textual state of the component after the modification.
Since the full text of a norm is constructed by hierarchically grouping the texts of its individual components, the creation of a new CTV for any single component automatically generates a new CTV for its parent components, all the way up to the Norm as a whole (for the norm as a whole, its is a TV). This ensures that the compiled (unofficial consolidated) text of the entire legal norm is always consistent at any given date.
For example, Figure 5 illustrates the Version entities for the provision (Component) “Art. 6º” (and its caput - main section) of ’Constituição Federal brasileira’ (Brazilian Federal Constitution)" [^10] The structuring of the ”Constituição Federal Brasileira de 1988” (Brazilian Federal Constitution of 1988) into ”Norm”, ”Components”, and ”Versions” can be visualized at https://normas.leg.br/?urn=urn:lex:br:federal:constituicao:1988-10-05;1988. Accessed on 15.03.2025.. In this monolingual example, each depicted "Version" entity represents both a Temporal Version and its corresponding Portuguese Language Version. The figure demonstrates how different versions, triggered by constitutional amendments, reflect changes over time.
Figure 5: Structure of the provision (Component) "Art. 6º" (and its caput - main section) of Brazilian Federal Constitution, illustrating four Version entities over time (light gray circles). Each represents a distinct Temporal Version (and its corresponding Portuguese Language Version) of the component, capturing the different wordings on different dates due to constitutional amendments.
@@ 2.5 Building the Graph
After semantically segmenting the text of a legal norm into its hierarchical structure, the elements of this hierarchy are inserted into the Knowledge Graph as nodes. As previously defined, we create nodes of type Norm (the Work) and Component (the Component Work), as illustrated in Figure 4. These nodes represent the conceptual backbone of the legal instrument, identified by properties like URN, name, and creation date.
The textual content itself is then linked to this abstract structure through Temporal Version and Language Version nodes. For each Component, we create:
  1. 1.
    A Temporal Version (CTV) node for each point in time its text was enacted or amended. This node is identified by a date-stamped URN.
  2. 2.
    A Language Version (CLV) node for each language in which that CTV’s text is available. This node is identified by a language-code-stamped URN.
Crucially, the text segment (text chunk) is associated with the Language Version node. This associated text is fundamental for the Retrieval-Augmented Generation process, as it represents the specific, contextualized content that will be retrieved and used by the LLM [14]. Figure 6 illustrates this multi-layered structure.
Figure 6: Representation of the multi-layered relationship in the graph: a Norm has Components, which have date-stamped Temporal Versions (CTVs), which in turn have language-specific Language Versions (CLVs). The Text Chunks are linked to the CLVs.
For example, the Norm (Work) node/entity “Constituição Federal Brasileira de 1988” (Brazilian Federal Constitution of 1988), identified by urn: [^11] The formation of the urns presented in this work follows the standard proposed by the Brazilian LexML project. Available at: https://projeto.lexml.gov.br. Accessed on: 15 Mar 2025.
It has a Component (Component Work) node/entity “caput do Art. 6º” (caput is the main text of the article) identified by:
This component has a original Temporal Version (Expression)—the textual formulation of the component on the date of norm’s signature:1988-10-05—is identified by adding @YYYY-MM-DD:
And this TV has a primary Language Version (Expression) for Portuguese ("pt"), identified by:
This Language Version node is associated with the Text Chunk containing the original wording (in Portuguese) of the caput of Art. 6º, as illustrated in Figure 5:
@@
Unlike static documents, the text of a legal component evolves. This evolution is represented by a set of Temporal Versions, each capturing a state of the text at a specific time. Figure 7 displays the four Temporal Versions of the caput of Art. 6º, with each CTV linked to its corresponding Portuguese CLV and text chunk—are displayed: one original CTV and three resulting from changes (text change) promoted by “Emendas Constitucionais” (Constitutional Amendments).
For a more precise representation in the Graph, specialized entity types can be derived from Temporal Version to represent the nature of the change found in the treated legal system. For example, in Brazil, there are types like "Original Version”, “Alteração de Texto” (Text Change Version), “Renumeração” (Renumbering Version), “Revogação” (Repeal Version) etc. Thus, the entity representing the temporal version could be included in the graph using the specialized type.
Figure 7: Representation of multiple Temporal Versions for the component (…!art6_cpt). Each CTV has a corresponding Language Version (in Portuguese) which is linked to its unique Text Chunk, illustrating the component’s textual evolution over time.
The core proposal of this work is to build a knowledge graph that models this complete hierarchical, temporal, and linguistic structure. It is important to note that text segments (text chunks) are linked to Language Version entities, not to Component or Temporal Version entities. This ensures that retrieval is always precise to a specific text in a specific language at a specific time.
This detailed representation allows answering questions about the text of a component with full temporal context: for example, if there was any change in the text over time; what changes it underwent; or what the text was on a specific date, and in which language.
@@ 3.1 Text Units
Regarding the text segment (text chunk) associated with each Language Version (LV) entity, in cases of subdivisions of a caput, such as “parágrafos” (paragraphs) and enumeration elements (“incisos” (items) and “alíneas” (subitems)), we suggest adopting the strategy proposed by [15]. In this strategy, the text segment to be considered for a component is not only the text directly resulting from the semantic segmentation of the norm but the semantically complete text of the component, derived from post-segmentation (and post-compilation) processing. The cited author recognizes the hierarchical and interdependent nature of legal norm components, where, for example, the meaning of an enumeration component (e.g., “inciso” (item)) is frequently dependent on the context of its parent component (e.g., the caput).
In other words, when including in the graph the original Language Version of the provision "inciso I do caput do Art. 5º” (item I of caput of Art. 5), the text linked to this entity will not be just:
@@
but rather the text formed by the content of the article’s caput’s CLV (from the same Temporal Version) plus the content of the item itself. This can be visualized in Figure 8, where the parent’s text was placed as a prefix in square brackets.
@@
Figure 8: Example of contextual (cumulative) Text Units for the Language Versions of the "incisos’ (items) of caput of Art. 5, demonstrating the inclusion of the parental caput text to provide complete semantic context.
Conversely, the Language Version of an article’s caput will have as its linked text its own text accumulated with the text from the CLVs of its child components (all belonging to the same Temporal Version), as exemplified in Figure 9. The same logic applies recursively up the hierarchy to articles, groupers, and the norm as a whole.
According to the cited author, this would bring more “aboutness” [^12] The term ”aboutness” refers to the capacity of a text to faithfully represent its central theme. For theoretical discussions, see [17]. to the texts linked to the entities [16]. In other words, the cumulative “Text Unit”, by including the textual context of parent or child components, represents the central theme and legal meaning of the component in question more completely and comprehensively.
Figure 9: Example of a cumulative Text Unit for the Language Version of the caput of Art. 5 of the "Constituição Federal Brasileira de 1988" (Brazilian Federal Constitution of 1988).
Considering this new approach, we will call the texts associated with Language Version entities Text Units, instead of Text Chunks. They are not necessarily the pieces of text resulting directly from the initial segmentation of the norms’ texts, but rather the contextually complete content of each component, ready for embedding and retrieval.
3.2 Temporal Versions as Aggregations
When a subsequent norm, on date D1, commands an alteration to the content of a component C, a new Temporal Version (CTV) is created for that component. This is illustrated in Figure 10, which shows a component’s lifecycle as a series of date-stamped CTVs.
Figure 10: Temporal Versions of component C over time. Each node represents a distinct textual state of the component on a specific date.
Considering the proposal of Text Units resulting from the accumulation of text in the hierarchy, consequently, a new Temporal Version CTV (and its related CLV) must also be produced for every component of norm N that: